OpenStudy (anonymous):

Write the equation in slope-intercept form for a line that passes through the points (-6, 0) and (0, -2).

OpenStudy (watchmath):

Give it a try! :D

OpenStudy (anonymous):

y-0=3(x-(-6))

OpenStudy (watchmath):

Be careful! the slope is -2/6 = -1/3.